There are five children in a family of parents AB× BB. The children of such parents must
have genotype AB or genotype BB . Find the probability that two of the children have
genotype AB and three others have genotype BB .
1
Expert's answer
2019-10-27T13:53:09-0400

Let probability of a child having genotype AB: "p=1\/2"

Then probability of a child having genotype BB: "1-p=1-1\/2=1\/2"

Let random variable "X=" number of children having genotype AB: "X\sim B(n, p)"


"P(X=x)=\binom{n}{x}p^x(1-p)^{n-x}"

Given that "n=5, p=1\/2"

 Find the probability that two of the children have genotype AB and three others have genotype BB .


"P(X=2)=\binom{5}{2}({1 \over 2})^2(1-{1 \over 2})^{5-2}={5 \over 16}=0.3125"